# The physical and chemical processes in protoplanetary disks: constraints on the composition of comets

### v3

@@ 1 INTRODUCTION
Before proceeding to the main sections, we briefly explain the process of star formation in order to set up the stage and to clarify the relationship with the chapter by Bergin et al. Low-mass stars like the Sun are formed by the gravitational collapse of a molecular cloud core. Since the core has non-zero angular momentum as a whole, the central star and its circumstellar disk are formed simultaneously. In the early stages, the star and disk are deeply embedded in dense gas, which we call envelope. Since the envelope gas is cold, the spectral energy distribution (SED) of the object has a peak in the far infrared. The objects in this evolutionary stage are called Class 0. In spite of its low temperature at the outermost radius of the envelope, the temperature of the forming disk would be high due to the high mass accretion rate inside the disk (i.e. release of gravitational energy), and shock heating caused by accretion from the envelope to the disk (e.g. Offner and McKee, 2011, see also Fig. 1 and §2.2). While the envelope gas accretes to the central star via disk, the high-velocity jet and outflow are launched from the vicinity of the central star. Eventually, the envelope gas decreases as it falls to the disk and star, or it is swept by the jet and outflow. When the central star (i.e. in near-infrared wavelengths) becomes apparent in the SED, which is still dominated by mid- to far-infrared though, the objects are called Class I. Eventually, when the envelope gas is dissipated, the central star becomes a T Tauri star and the disk becomes a protoplanetary disk, which is also called Class II disk. More detailed explanations and quantitative definitions of Class 0, I, and II are found in Evans et al. (2009).
@@
While it is not easy to determine the age of pre-main-sequence stars, the typical age of the Class II objects is a few years. Statistical observations (i.e. number counts) indicate that the lifetime for Class 0 and Class I is Myr and Myr, respectively (e.g. White et al., 2007; Evans et al., 2009). In order to set Solar system formation in an astronomical context, these timescales are compared with the range of absolute ages (i.e. formation interval) of Calcium and Aluminium Inclusions (CAIs) and chondrules in meteorites. CAIs have the highest condensation temperature among minerals in chondrites and are thus considered to be formed in the hottest stage and/or region. The formation interval of CAIs is estimated to be 0.16 Myr by the isotope dating, which is comparable to the lifetime of Class 0 stage. The age of the chondrules, on the other hand, varies over a few Myr, which coincides with the lifetime of Class II disks (Connelly et al., 2012). While Class II disks have been considered to be the birthplace of planetary systems, a ring-gap structure of dust thermal emission is recently found in young disks in Class 0 to I as well. It suggests that planetary system formation may start early (see §4). In summary, the Solar nebula may correspond to the whole evolutionary stages from Class 0, I, and to II.
@@ 2.1 Gas dynamics
From the observational side, direct constraints on disk turbulence strength have been obtained from measurements of nonthermal Doppler broadening of molecular emission lines (Hughes et al., 2011; Guilloteau et al., 2012). Recently, there have been several attempts to detect disks’ nonthermal gas motions using ALMA, but most of them have resulted in non-detection with upper limits on the nonthermal velocity dispersion of several to ten % of the local sound speed (Flaherty et al., 2015, 2017, 2018; Teague et al., 2016, 2018). To date, the disk around DM Tau is the only case for which turbulent motion has been detected with ALMA (Flaherty et al., 2020), albeit at a few gas scale heights above the midplane. There are indirect constraints on turbulence strength at the midplane from the morphology or azimuthal emission variation of dust rings and gaps seen in ALMA millimeter continuum images (Pinte et al. 2016; Dullemond et al. 2018; Rosotti et al. 2020; Doi and Kataoka 2021; see Sect. 2.5 for more detail about disk substructures found by millimeter observations). Some rings show an indication of strong vertical settling or radial dust concentration, suggesting that the turbulence at the midplane of the disks is too weak to diffuse the dust particles. However, constraining the level of turbulence quantitatively from these observations requires additional constraints on the dust particle size because large dust particles may settle and concentrate even in the presence of strong turbulence.
@@ 2.2 Temperature distribution and snowlines
Themodynamically, the snowline of each volatile species can be defined as the location where its saturation vapor pressure equals its partial pressure in the gas phase. To illustrate how the locations of the snowlines are determined, we plot in Fig. 2 the saturation vapor pressure curves for pure , , , and ices as a function of radial distance for the classical, optically thin minimum-mass Solar nebula (MMSN) model of Hayashi (1981), which has radial temperature and midplane pressure profiles of and , respectively. Here, we adopt analytic expressions for the saturation vapor pressures as a function of temperature provided by Bauer et al. (1997) (for HO) and by Yamamoto et al. (1983) (for the other species), and plot them as a function of using the assumed temperature profile. Let us assume that the disk contains , , , and at spatially uniform molar abundances of , , , and , respectively, so that the relative abundances between the four volatiles are crudely consistent with the abundances in comets (Mumma and Charnley, 2011). If all the volatiles were in the gas phase, they would have partial pressures of . By comparing the saturation vapor pressures and assumed partial pressures, we find that the snowlines of , , , and are located at 206, 17, 12, and 3.2 au, with sublimation temperatures of , , , and K, respectively, in this particular disk model (see the circles in Fig. 2). [The abundance of HO relative to H is set to assuming that HO is the dominant reservoir of oxygen (Lodders, 2003). If water is mostly inherited from molecular clouds, HO abundance would be (Whittet, 1993). Then the water snowline is slightly shifted outwards. We also note that the ices would be a mixture of various molecules, while we consider pure ices here for simplicity. Sublimation of mixed ice is more complicated than that of pure ice (e.g. see review by Hama and Watanabe, 2013; Minissale et al., 2022). For example, laboratory experiments show that the adsorption energy and thus the sublimation temperature of a molecule depends on the composition and surface structure (e.g. crystal or amorphous) of the substrate ice. Referring to the observations of comets and interstellar ice, HO is expected to be a dominant composition of ice in disks. Then a fraction of molecule with higher volatility, such as CO and CO, would be trapped in HO ice and desorb at HO snow line.]
@@
The large temperature difference between the midplane and surface regions of the disks has significant implications for the compositional distribution of solids in the disks. Fig. 3(a) illustrates how the locations of the snowlines of some major volatiles at the surface (here referring to three scale heights above the midplane) and at the midplane evolve in an optically thick, passively irradiated disk (see also the snow surface in Fig. 1). Here, the central star is assumed to have a fixed mass of , while the stellar luminosity is assumed to decrease from at to at in a power-law fashion, mimicking the evolution of solar-mass pre-main-sequence stars. For simplicity, the abundances of the volatiles are assumed to be spatially uniform and equal to the values used in Fig. 2, and the surface density profile of the MMSN is used to calculate the disk pressure distribution. We find from the lower panel of Fig. 3 (a) that the HO snowline at the midplane would move inside 1 au—the radius of the current Earth’s orbit—within the first million years of star and planet formation. In contrast, the HO snowline in the optically thin surface region lies farther away, at more than au from the central star (see the upper panel of Fig. 3 (a)). This difference arises mainly from the vertical temperature (rather than pressure) gradient.
The midplane temperature given by Equation (2) should be taken as a lower limit because we have neglected any internal heat sources. The classical accretion disk model that assumes vertically uniform turbulent viscosity (Lynden-Bell and Pringle, 1974) predicts that the gravitational energy liberated by mass accretion toward the star can significantly heat the disk interior, in particular in the early disk evolutionary phase where both the accretion rate and disk surface density are high. Fig. 3 (b) illustrates the evolution of the midplane temperature for a disk heated by internal viscous heating in addition to stellar radiation. We consider a solar-mass star and let its luminosity evolve as in Fig. 3 (a). The disk accretion rate is assumed to decay as (Hartmann et al., 2016). The disk opacity is taken to be (Bell and Lin, 1994), and the disk’s turbulent viscosity is taken to be 1 % of the local sound velocity times the gas scale height. The figure shows that the viscous heating pushes the HO snowline out beyond 1 au during the entire Class II phase ( several Myr). Note that the temperature evolution illustrated here depends on the disk opacity assumed, and hence on the evolution of dust grains that dominate the opacity. In the disk model adopted here, a factor of 10 decrease in the opacity leads to a factor of two decrease in the midplane temperature.
@@
The disk temperature structure and the locations of the snowlines drastically change during episodic accretion outburst events introduced in Sect. 2.1. For example, V883 Ori is a solar-mass FU Ori star with an elevated luminosity of (Furlan et al., 2016) and a massive circumstellar disk. Equation (2) implies that the midplane of the V883 Ori disk in the outburst phase should be times hotter than those of disks around solar-luminosity stars. Assuming a sublimation temperature of 150 K for HO ice, we find that the strong irradiation heating should push the HO snowline out to 20 au. In reality, the bolometric luminosity of includes accretion heating, which could push the snowline further out by heating the disk from inside.
@@ 2.3 Dust growth: theory
In the context of this chapter, it is important to point out that grains’ chemical composition is key to understanding how they grow. Early models and experiments suggested that aggregates made of water ice are sticker than silicates (Dominik and Tielens, 1997) and may even overcome the fragmentation barrier (Wada et al., 2009). Such sticky grains also tend to form highly porous aggregates, which is beneficial for overcoming the drift barrier because the porous aggregates grow rapidly (Okuzumi et al., 2012; Kataoka et al., 2013). However, this sticky water ice scenario is questioned by recent experiments showing that water ice is not so sticky at low temperatures (Gundlach et al., 2018; Musiolik and Wurm, 2019). Instead, some recent studies suggest that silicates are sticker than previously thought (Kimura et al., 2015; Steinpilz et al., 2019). Some (if not all) types of organic matter are also sticky in a warm environment (Kouchi et al., 2002; Piani et al., 2017; Bischoff et al., 2020). CO ice appears to be less sticky than water ice (Musiolik et al., 2016a, b; Arakawa and Krijt, 2021; Fritscher and Teiser, 2021). These imply that the fate of dust coagulation and planetesimal formation may depend on temperature and hence on distance from the central star (Birnstiel et al., 2010; Pinilla et al., 2017; Homma et al., 2019; Okuzumi and Tazaki, 2019).
@@ 2.4 Dust growth: observations
Recently, radio polarimetric observations have been used to derive independent constraints on the grain size in the outer regions ( 10 of au) of protoplanetary disks. The observations have shown that many inclined disks produce uniformly polarized submillimeter continuum emission whose polarization direction is parallel to the disks’ minor axes (e.g., Stephens et al., 2014, 2017; Hull et al., 2018; Harris et al., 2018; Cox et al., 2018; Sadavoy et al., 2018; Dent et al., 2019). This polarization pattern can be explained if the polarized continuum emission is dust thermal radiation scattered by the dust particles themselves (Kataoka et al., 2015; Yang et al., 2016). In this interpretation, the observed degree of polarization gives a strong constraint on the size of the opacity-dominating dust particles because the scattering and polarization efficiencies depend strongly on grain size. Models assuming spherical, compact silicate particles (Kataoka et al., 2015; Yang et al., 2016) predict a maximum grain size of for the disks with uniformly polarized submillimeter emission.
@@ 2.5 Dust spatial distribution
In principle, snowlines could also produce dust rings and gaps as mentioned in Section 2.3. Zhang et al. (2015) and Okuzumi et al. (2016) argued that the snowlines of major volatiles, including HO and CO, are responsible for the major dust rings and gaps in the HL Tau disk. However, the current understanding is that most of the rings and gaps that have been observed in a number of disks are not relevant to snowlines (Huang et al., 2018; Long et al., 2018; van der Marel et al., 2019). The main reason is that the radial positions of the substructures do not appear to be correlated with the central star’s luminosity, which should determine snowline locations. Nevertheless, there are a few disks (HD 163296 and MWC 480 in addition to HL Tau) in which a dust ring is found close to the expected location of the CO snowline (Zhang et al., 2021).
ALMA observations of dust emission have also provided important constraints on the dust surface density in the outer () disk region where icy bodies like comets form. If we assume that this region is optically thin to its own millimeter emission, the dust surface density is derived from the millimeter intensity as (see also Section 2.4). For example, the dust surface densities of the IM Lup, GM Aur, and AS 209 disks are estimated to be comparable or an order of magnitude higher than that in the MMSN ((ice+rock) g cm) for a reasonable assumption about for 0.1–1 mm sized grains (see the bottom panels of Fig. 5 in Section 3). We note, however, that these particular objects are among the largest, most massive disks; their dust surface densities may not represent those of more common, compact disks (Miotello et al., 2023).
@@ 3 Chemistry in Protplanetary disks
Comets contain a significant amount of icy material, which has been investigated to reveal the thermal history and chemical environment in the cold outer regions of the Solar nebula (e.g. chapter by Biver et al.). For example, the composition of major cometary ice is often similar to that of interstellar ice. The relative abundance of CO and CHOH to water is and in comets, while these ratios are and in quiescent molecular clouds (Mumma and Charnley, 2011). Inheritance of interstellar ice in comets has thus long been discussed. Observations of protoplanetary disks, on the other hand, can reveal the spatial distribution and evolution of molecules at the sites of ongoing planetary-system formation. These two kinds of information are complementary to elucidate the formation of comets and the Solar system.
@@
Theoretical and observational studies show that various chemical processes are going on within the disks (e.g. Aikawa et al., 2002; Henning and Semenov, 2013; Öberg et al., 2023). As we have seen in the previous section, there are large gradients of temperature and density in both the radial and vertical directions in a disk (Fig. 1). In the vertical directions, gases are in hydrostatic equilibrium and the density decreases towards the disk surface. The disk surface is directly irradiated by the stellar radiation and thus is warm. Molecules are dissociated to radicals and atoms by UV radiation from the central star 111While the surface temperature of T Tauri stars is K, UV radiation is produced by the gas accretion from the disk to the star. and interstellar radiation field; the disk surface is a photon-dominated region (PDR). Beneath the surface layer, there is a warm molecular layer ( K), which is moderately shielded from UV radiation and thus harbors various molecular gas such as CO. Ion-molecule reactions are triggered by X-rays from the central object 222X-rays are emitted due to magnetic activity on the stellar surface. and cosmic rays. In the midplane, the density is highest and dust grains could be further concentrated due to sedimentation. Since the temperature decreases with radial distance, snowlines are defined in the midplane as explained in §2.2 (Fig. 2). Thermal energy could be high enough to drive chemical reactions in the inner hot regions (e.g. above several hundreds of K at au, Yang et al., 2013). Even in the outer cold midplane, reactions in the gas phase and within ices can be triggered by high-energy cosmic rays, X-rays, and decay of radioactive nuclei (e.g. Al). Such layered structures, i.e. the PDR layer, warm molecular layer, and icy cold midplane, are confirmed by spatially and spectrally resolved observations (i.e. channel maps) of disks (Dutrey et al., 2017; Ruíz-Rodríguez et al., 2021; Rosenfeld et al., 2013; Law et al., 2021b). While the chemical composition of ice in the midplane would be most directly relevant to comets, observation of such ice is not easy (see §3.3). Gaseous line observations in millimeter and submillimeter wavelengths are much more sensitive to low-abundance species. Chemistry in the gas and ice is linked via freeze-out, sublimation, and radial/vertical transport as we will discuss below.
@@
DNC | sub-mm | [18] | HCO | sub-mm, mm | [8] [29] |
DCO | sub-mm, mm | [30] [27] | HCO | sub-mm, mm | [30] [25] |
HCO | sub-mm | [31] | NH | mm | [32] [33] |
ND | mm | [34] [35] | HS | mm | [36] [37] |
CO | MIR | [16][56] | CO | MIR | [56] |
CS | mm | [53] | SO | sub-mm | [38] |
SO | sub-mm | [38] | |||
NH | MIR, sub-mm | [39][58] | CH | NIR, MIR | [40] [56] |
CCH | MIR | [57] | HCO | sub-mm, mm | [8] [41] |
HCS | sub-mm, mm | [17] [18] | HCN | mm | [42] [43] |
HCOOH | mm | [44] | CH | mm | [45] [43] |
CH | NIR | [46] | CHCN | mm | [47] |
CHCN | mm | [48] [43] | CHOH | sub-mm, mm | [49] [50] |
CH | MIR | [57] | CHCHO | submm | [52] |
CHOCHO | submm | [51] [52] | CHOCH | submm | [51] |
CHCOCH | submm | [52] | CH | MIR | [57] |
aAdopted and revised from McGuire (2018). The range of wavelengths are 2-5 m for near-indrared (NIR), 5-40 m for mid-infrared (MIR), 40-300 m for far-infrared (FIR), 300-1000 m for sub-millimeter (sub-mm), and mm for millimeter (mm). References are not exhaustive. In addition to the first detection papers, we list recent papers for readers to refer to references therein (see also text).
@@
[57] Tabone et al. (2023) [58] Najita et al. (2021)

### v4

@@ 1 INTRODUCTION
Before proceeding to the main sections, we briefly explain the process of star formation in order to set up the stage and to clarify the relationship with the chapter by Bergin et al. Low-mass stars like the Sun are formed by the gravitational collapse of a molecular cloud core. Since the core has non-zero angular momentum as a whole, the central star and its circumstellar disk are formed simultaneously. In the early stages, the star and disk are deeply embedded in dense gas, which we call envelope. Since the envelope is cold, the spectral energy distribution (SED) of the object has a peak in the far infrared. The objects in this evolutionary stage are called Class 0. In spite of its low temperature at the outermost radius of the envelope, the temperature of the forming disk would be high due to the high mass accretion rate inside the disk (i.e. release of gravitational energy), and shock heating caused by accretion from the envelope to the disk (e.g. Offner and McKee, 2011, see also Fig. 1 and §2.2). While the envelope gas accretes to the central star via disk, the high-velocity jet and outflow are launched from the vicinity of the central star. Eventually, the envelope gas decreases as it falls to the disk and star, or it is swept by the jet and outflow. When the central star (i.e. in near-infrared wavelengths) becomes apparent in the SED, which is still dominated by mid- to far-infrared though, the objects are called Class I. Eventually, when the envelope gas is dissipated, the central star becomes a T Tauri star and the disk becomes a protoplanetary disk, which is also called Class II disk. More detailed explanations and quantitative definitions of Class 0, I, and II are found in Evans et al. (2009).
@@
While it is not easy to determine the age of pre-main-sequence stars, the typical age of the Class II objects is a few years. Statistical observations (i.e. number counts) indicate that the lifetime for Class 0 and Class I is Myr and Myr, respectively (e.g. White et al., 2007; Evans et al., 2009). In order to set Solar system formation in an astronomical context, these timescales are compared with the range of absolute ages (i.e. formation interval) of Calcium and Aluminium Inclusions (CAIs) and chondrules in meteorites. CAIs have the highest condensation temperature among minerals in chondrites and are thus considered to be formed in the hottest stage and/or region. The formation interval of CAIs is estimated to be 0.16 Myr by the isotope dating, which is comparable to the lifetime of Class 0 stage. The age of the chondrules, on the other hand, varies over a few Myr, which coincides with the lifetime of Class II disks (Connelly et al., 2012). While Class II disks have been considered to be the birthplace of planetary systems, a ring-gap structure of dust thermal emission is recently found in some Class I disks as well. It suggests that planetary system formation may start early (see §4). In summary, the Solar nebula may correspond to the whole evolutionary stages from Class 0, I, and to II.
@@ 2.1 Gas dynamics
From the observational side, direct constraints on disk turbulence strength have been obtained from measurements of nonthermal Doppler broadening of molecular emission lines (Hughes et al., 2011; Guilloteau et al., 2012). Recently, there have been several attempts to detect disks’ nonthermal gas motions using ALMA, but most of them have resulted in non-detection with upper limits on the nonthermal velocity dispersion of several to ten % of the local sound speed (Flaherty et al., 2015, 2017, 2018; Teague et al., 2016, 2018). To date, the disks around DM Tau (Flaherty et al., 2020) and IM Lup (Paneque-Carreño et al., 2024; Flaherty et al., 2024) are the only cases for which turbulent motion has been detected with ALMA , albeit at a few gas scale heights above the midplane. There are indirect constraints on turbulence strength at the midplane from the morphology or azimuthal emission variation of dust rings and gaps seen in ALMA millimeter continuum images (Pinte et al. 2016; Dullemond et al. 2018; Rosotti et al. 2020; Doi and Kataoka 2021; see Sect. 2.5 for more detail about disk substructures found by millimeter observations). Some rings show an indication of strong vertical settling or radial dust concentration, suggesting that the turbulence at the midplane of the disks is too weak to diffuse the dust particles. However, constraining the level of turbulence quantitatively from these observations requires additional constraints on the dust particle size because large dust particles may settle and concentrate even in the presence of strong turbulence.
@@ 2.2 Temperature distribution and snowlines
Themodynamically, the snowline of each volatile species can be defined as the location where its saturation vapor pressure equals its partial pressure in the gas phase. To illustrate how the locations of the snowlines are determined, we plot in Fig. 2 the saturation vapor pressure curves for pure , , , and ices as a function of radial distance for the classical, optically thin minimum-mass Solar nebula (MMSN) model of Hayashi (1981), which has radial temperature and midplane pressure profiles of and , respectively. Here, we adopt analytic expressions for the saturation vapor pressures as a function of temperature provided by Bauer et al. (1997) (for H2O) and by Yamamoto et al. (1983) (for the other species), and plot them as a function of using the assumed temperature profile. Let us assume that the disk contains , , , and at spatially uniform molar abundances of , , , and , respectively, so that the relative abundances between the four volatiles are crudely consistent with the abundances in comets (Mumma and Charnley, 2011). If all the volatiles were in the gas phase, they would have partial pressures of . By comparing the saturation vapor pressures and assumed partial pressures, we find that the snowlines of , , , and are located at 206, 17, 12, and 3.2 au, with sublimation temperatures of , , , and K, respectively, in this particular disk model (see the circles in Fig. 2). [The abundance of H2O relative to H2 is set to assuming that H2O is the dominant reservoir of oxygen (Lodders, 2003). If water is mostly inherited from molecular clouds, H2O abundance would be (Whittet, 1993). Then the water snowline is slightly shifted outwards. We also note that the ices would be a mixture of various molecules, while we consider pure ices here for simplicity. Sublimation of mixed ice is more complicated than that of pure ice (e.g. see review by Hama and Watanabe, 2013; Minissale et al., 2022). For example, laboratory experiments show that the adsorption energy and thus the sublimation temperature of a molecule depends on the composition and surface structure (e.g. crystal or amorphous) of the substrate ice. Referring to the observations of comets and interstellar ice, H2O is expected to be a dominant composition of ice in disks. Then a fraction of molecule with higher volatility, such as CO and CO2, would be trapped in H2O ice and desorb at H2O snow line.]
@@
The large temperature difference between the midplane and surface regions of the disks has significant implications for the compositional distribution of solids in the disks. Fig. 3(a) illustrates how the locations of the snowlines of some major volatiles at the surface (here referring to three scale heights above the midplane) and at the midplane evolve in an optically thick, passively irradiated disk (see also the snow surface in Fig. 1). Here, the central star is assumed to have a fixed mass of , while the stellar luminosity is assumed to decrease from at to at in a power-law fashion, mimicking the evolution of solar-mass pre-main-sequence stars. For simplicity, the abundances of the volatiles are assumed to be spatially uniform and equal to the values used in Fig. 2, and the surface density profile of the MMSN is used to calculate the disk pressure distribution. We find from the lower panel of Fig. 3 (a) that the H2O snowline at the midplane would move inside 1 au—the radius of the current Earth’s orbit—within the first million years of star and planet formation. In contrast, the H2O snowline in the optically thin surface region lies farther away, at more than au from the central star (see the upper panel of Fig. 3 (a)). This difference arises mainly from the vertical temperature (rather than pressure) gradient.
The midplane temperature given by Equation (2) should be taken as a lower limit because we have neglected any internal heat sources. The classical accretion disk model that assumes vertically uniform turbulent viscosity (Lynden-Bell and Pringle, 1974) predicts that the gravitational energy liberated by mass accretion toward the star can significantly heat the disk interior, in particular in the early disk evolutionary phase where both the accretion rate and disk surface density are high. Fig. 3 (b) illustrates the evolution of the midplane temperature for a disk heated by internal viscous heating in addition to stellar radiation. We consider a solar-mass star and let its luminosity evolve as in Fig. 3 (a). The disk accretion rate is assumed to decay as (Hartmann et al., 2016). The disk opacity is taken to be (Bell and Lin, 1994), and the disk’s turbulent viscosity is taken to be 1 % of the local sound velocity times the gas scale height. The figure shows that the viscous heating pushes the H2O snowline out beyond 1 au during the entire Class II phase ( several Myr). Note that the temperature evolution illustrated here depends on the disk opacity assumed, and hence on the evolution of dust grains that dominate the opacity. In the disk model adopted here, a factor of 10 decrease in the opacity leads to a factor of two decrease in the midplane temperature.
@@
The disk temperature structure and the locations of the snowlines drastically change during episodic accretion outburst events introduced in Sect. 2.1. For example, V883 Ori is a solar-mass FU Ori star with an elevated luminosity of (Furlan et al., 2016) and a massive circumstellar disk. Equation (2) implies that the midplane of the V883 Ori disk in the outburst phase should be times hotter than those of disks around solar-luminosity stars. Assuming a sublimation temperature of 150 K for H2O ice, we find that the strong irradiation heating should push the H2O snowline out to 20 au. In reality, the bolometric luminosity of includes accretion heating, which could push the snowline further out by heating the disk from inside.
@@ 2.3 Dust growth: theory
In the context of this chapter, it is important to point out that grains’ chemical composition is key to understanding how they grow. Early models and experiments suggested that aggregates made of water ice are sticker than silicates (Dominik and Tielens, 1997) and may even overcome the fragmentation barrier (Wada et al., 2009). Such sticky grains also tend to form highly porous aggregates, which is beneficial for overcoming the drift barrier because the porous aggregates grow rapidly (Okuzumi et al., 2012; Kataoka et al., 2013). However, this sticky water ice scenario is questioned by recent experiments showing that water ice is not so sticky at low temperatures (Gundlach et al., 2018; Musiolik and Wurm, 2019). Instead, some recent studies suggest that silicates are sticker than previously thought (Kimura et al., 2015; Steinpilz et al., 2019). Some (if not all) types of organic matter are also sticky in a warm environment (Kouchi et al., 2002; Piani et al., 2017; Bischoff et al., 2020). CO2 ice appears to be less sticky than water ice (Musiolik et al., 2016a, b; Arakawa and Krijt, 2021; Fritscher and Teiser, 2021). These imply that the fate of dust coagulation and planetesimal formation may depend on temperature and hence on distance from the central star (Birnstiel et al., 2010; Pinilla et al., 2017; Homma et al., 2019; Okuzumi and Tazaki, 2019).
@@ 2.4 Dust growth: observations
Recently, radio polarimetric observations have been used to derive independent constraints on the grain size in the outer regions ( 10 au) of protoplanetary disks. The observations have shown that many inclined disks produce uniformly polarized submillimeter continuum emission whose polarization direction is parallel to the disks’ minor axes (e.g., Stephens et al., 2014, 2017; Hull et al., 2018; Harris et al., 2018; Cox et al., 2018; Sadavoy et al., 2018; Dent et al., 2019). This polarization pattern can be explained if the polarized continuum emission is dust thermal radiation scattered by the dust particles themselves (Kataoka et al., 2015; Yang et al., 2016). In this interpretation, the observed degree of polarization gives a strong constraint on the size of the opacity-dominating dust particles because the scattering and polarization efficiencies depend strongly on grain size. Models assuming spherical, compact silicate particles (Kataoka et al., 2015; Yang et al., 2016) predict a maximum grain size of for the disks with uniformly polarized submillimeter emission.
@@ 2.5 Dust spatial distribution
In principle, snowlines could also produce dust rings and gaps as mentioned in Section 2.3. Zhang et al. (2015) and Okuzumi et al. (2016) argued that the snowlines of major volatiles, including H2O and CO, are responsible for the major dust rings and gaps in the HL Tau disk. However, the current understanding is that most of the rings and gaps that have been observed in a number of disks are not relevant to snowlines (Huang et al., 2018; Long et al., 2018; van der Marel et al., 2019). The main reason is that the radial positions of the substructures do not appear to be correlated with the central star’s luminosity, which should determine snowline locations. Nevertheless, there are a few disks (HD 163296 and MWC 480 in addition to HL Tau) in which a dust ring is found close to the expected location of the CO snowline (Zhang et al., 2021).
ALMA observations of dust emission have also provided important constraints on the dust surface density in the outer () disk region where icy bodies like comets form. If we assume that this region is optically thin to its own millimeter emission, the dust surface density is derived from the millimeter intensity as (see also Section 2.4). For example, the dust surface densities of the IM Lup, GM Aur, and AS 209 disks are estimated to be comparable or an order of magnitude higher than that in the MMSN ((ice+rock) g cm-2) for a reasonable assumption about for 0.1–1 mm sized grains (see the bottom panels of Fig. 5 in Section 3). We note, however, that these particular objects are among the largest, most massive disks; their dust surface densities may not represent those of more common, compact disks (Miotello et al., 2023).
@@ 3 Chemistry in Protplanetary disks
Comets contain a significant amount of icy material, which has been investigated to reveal the thermal history and chemical environment in the cold outer regions of the Solar nebula (e.g. chapter by Biver et al.). For example, the composition of major cometary ice is often similar to that of interstellar ice. The relative abundance of CO2 and CH3OH to water is and in comets, while these ratios are and in quiescent molecular clouds (Mumma and Charnley, 2011). Inheritance of interstellar ice in comets has thus long been discussed. Observations of protoplanetary disks, on the other hand, can reveal the spatial distribution and evolution of molecules at the sites of ongoing planetary-system formation. These two kinds of information are complementary to elucidate the formation of comets and the Solar system.
@@
Theoretical and observational studies show that various chemical processes are going on within the disks (e.g. Aikawa et al., 2002; Henning and Semenov, 2013; Öberg et al., 2023). As we have seen in the previous section, there are large gradients of temperature and density in both the radial and vertical directions in a disk (Fig. 1). In the vertical directions, gases are in hydrostatic equilibrium and the density decreases towards the disk surface. The disk surface is directly irradiated by the stellar radiation and thus is warm. Molecules are dissociated to radicals and atoms by UV radiation from the central star 111While the surface temperature of T Tauri stars is K, UV radiation is produced by the gas accretion from the disk to the star. and interstellar radiation field; the disk surface is a photon-dominated region (PDR). Beneath the surface layer, there is a warm molecular layer ( K), which is moderately shielded from UV radiation and thus harbors various molecular gas such as CO. Ion-molecule reactions are triggered by X-rays from the central object 222X-rays are emitted due to magnetic activity on the stellar surface. and cosmic rays. In the midplane, the density is highest and dust grains could be further concentrated due to sedimentation. Since the temperature decreases with radial distance, snowlines are defined in the midplane as explained in §2.2 (Fig. 2). Thermal energy could be high enough to drive chemical reactions in the inner hot regions (e.g. above several hundreds of K at au, Yang et al., 2013). Even in the outer cold midplane, reactions in the gas phase and within ices can be triggered by high-energy cosmic rays, X-rays, and decay of radioactive nuclei (e.g. 26Al). Such layered structures, i.e. the PDR layer, warm molecular layer, and icy cold midplane, are confirmed by spatially and spectrally resolved observations (i.e. channel maps) of disks (Dutrey et al., 2017; Ruíz-Rodríguez et al., 2021; Rosenfeld et al., 2013; Law et al., 2021b). While the chemical composition of ice in the midplane would be most directly relevant to comets, observation of such ice is not easy (see §3.3). Gaseous line observations in millimeter and submillimeter wavelengths are much more sensitive to low-abundance species. Chemistry in the gas and ice is linked via freeze-out, sublimation, and radial/vertical transport as we will discuss below.
@@
| DNC | sub-mm | [18] | HCO+ | sub-mm, mm | [8] [29] |
| DCO+ | sub-mm, mm | [30] [27] | H13CO+ | sub-mm, mm | [30] [25] |
| HC18O+ | sub-mm | [31] | N2H+ | mm | [32] [33] |
| N2D+ | mm | [34] [35] | H2S | mm | [36] [37] |
| CO2 | MIR | [16][56] | 13CO2 | MIR | [56] |
| C2S | mm | [53] | OCS | mm | [60] |
| SO2 | sub-mm | [38] | 34SO2 | sub-mm | [38] |
| NH3 | MIR, sub-mm | [39][58] | C2H2 | NIR, MIR | [40] [56] |
| 13C12CH2 | MIR | [57] | H2CO | sub-mm, mm | [8] [41] |
| H2CS | sub-mm, mm | [17] [18] | HC3N | mm | [42] [43] |
| HCOOH | mm | [44] | C3H2 | mm | [45] [43] |
| CH4 | NIR | [46] | CH2CN | mm | [47] |
| CH3CN | mm | [48] [43] | CH3OH | sub-mm, mm | [49] [50] |
| C4H2 | MIR | [57] | CH3CHO | submm | [52] |
| CH3OCHO | submm | [51] [52] | 13CH3OCHO | mm | [60] |
| CH3O13CHO | mm | [60] | CH3OCH3 | submm | [51] [59] |
| CH3COCH3 | submm | [52] | C2H3CHO | mm | [60] |
| C2H5CHO | mm | [60] | C6H6 | MIR | [57] |
aAdopted and revised from McGuire (2018). This table includes molecules detected in the disk of V883 Ori, which is a Class I object in outburst phase (section 4.2). The range of wavelengths are 2-5 m for near-indrared (NIR), 5-40 m for mid-infrared (MIR), 40-300 m for far-infrared (FIR), 300-1000 m for sub-millimeter (sub-mm), and mm for millimeter (mm). References are not exhaustive. In addition to the first detection papers, we list recent papers for readers to refer to references therein (see also text).
@@
[57] Tabone et al. (2023) [58] Najita et al. (2021) [59] Yamato et al. (2024a)
[60] Yamato et al. (2024b)